OpenStudy (anonymous):

Show and Explain your work for this problem. (-4)(-3)/6-2(5-(-8)+(6/2))=

OpenStudy (ghazi):

12/6-2(5+8+6/2)=2-2(5+8+3)= 2-2(16)=2-32=-30

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:)